OpenStudy (anonymous):

Ellie had 44 books in here bookcase. She counted 32 fiction book and the rest were nonfiction. What was the ratio of fiction books to nonfiction book in Ellie's bookcase? A. 8 to 3 B. 8 to 11 C. 11 to 3 D. 3 to 8

OpenStudy (anonymous):

@Lyralei

OpenStudy (anonymous):

First let's find the number of nonfiction books. So if there are 32 fiction and 44 total, to find the nonfiction just do 44-32.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Right. So now we put together our ratio. Keep in mind that \[\large x~to~y~=~\frac{x}{y}\] so \[\large fiction~to~nonfiction~=~\frac{fiction}{nonfiction}\]

OpenStudy (anonymous):

So if we substitute the values of fiction and nonfiction in, what would it look like?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

Well how many fiction books are there and how many nonfiction books are there?

OpenStudy (anonymous):

32 fiction and 12 nonfiction

OpenStudy (anonymous):

Right. So susbtitute the number of fiction and the number of nonfiction into \[\large \frac{fiction}{nonfiction}\]

OpenStudy (anonymous):

you mean like 12/32

OpenStudy (anonymous):

The opposite \[\large \frac{32}{12}\] but anyways now it's time to simplify. Do you know how to do that?

OpenStudy (anonymous):

8/3

OpenStudy (anonymous):

Right! So now, we just have to translate it back to the other ratio form. If \[\large \frac{x}{y}~=~x~to~y\] then what will \[\large \frac{8}{3}\]be?

OpenStudy (anonymous):

8 to 3

OpenStudy (anonymous):

Correct! Good job! :)

OpenStudy (anonymous):

Thank you for the help

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]